OpenStudy (anonymous):

x + y = k x - y = k The solution to the system shown is _____. (0, k) (2, 2) (k, 0)

OpenStudy (texaschic101):

x + y = k x - y = k ---------add 2x = 2k -- divide both sides by 2 x = k x + y = k k + y = k y = k - k y = 0 x + 0 = k x = k - 0 x = k solution is : (k,0)......I think

OpenStudy (anonymous):

@texaschic101 you are right

OpenStudy (anonymous):

Thank you! =))